# Physics-informed Mesh-independent Deep Compositional Operator Network

### v1

@@ 4.3 Comparison with data-driven neural operators
One of the main features of PI-DCON is that it is a data-free approach. It doesn’t require FE simulation runs as a training set. In this section, we seek to investigate its efficiency and accuracy in comparison with data-driven (supervised) training approaches. In doing so, we consider the data-driven DeepONet and also create a data-driven variant of our proposed DCON architecture for comparison. We first compare the performance of data-driven DCON with that of data-driven DeepONet in Table LABEL:table.data_driven_com. We observe that the DCON model achieved 44% accuracy improvement for the Darcy flow problem and 58% accuracy improvement for the 2D plate problem. Also, when we compare our model’s performance on data with variable mesh sizes to those with a fixed mesh size, we do not observe a significant drop in the prediction accuracy, which is comparable to the drop observed in PI-DCON.
@@
Dataset Mean of relative error Standard deviation of relative error
Data-driven DeepONet Data-driven DCON Data-driven DeepONet Data-driven DCON
Darcy flow (fixed mesh size) 3.74% 2.09% 0.97% 0.51%
2D plate (fixed mesh size) 3.78% 1.57% 1.01% 0.37%
Darcy flow (variable mesh size) - 2.22% - 0.59%
2D plate (variable mesh size) - 1.78% - 0.35%
In order to demonstrate the advantage of physics-informed (data-free) training of DCON, in Figure 7 we show the required timea need for building models at different accuracy levels. The horizontal (dashed) line represents the required CPU time for generating training datasets for data-driven training of neural operators. In both darcy flow and 2d plate problems, the lowest error rate achieved by PI-DCON is smaller than that of data-driven DeepONet but greater than that of data-driven DCON. This result suggests that neural operators trained in a data-driven fashion tend to outperform those trained using physics-informed approaches when the same model architecture is used. However, it can be seen that PI-DCON convergence is superior compared to data-driven DeepONet. This shows the enhanced architectural design in DCON improves alows for faster convergence compared to the conventional DeepONet.
@@ 4.4 Ablation studies
Darcy flow (fixed mesh size) 10.10% 4.20% 2.50 % 2.15% 2.06% 2.03%
2D plate (fixed mesh size) 5.92% 2.57% 1.77% 1.56% 1.43% 1.38%
Darcy flow (variable mesh size) 12.58% 5.01% 3.42% 2.83% 2.66% 2.59%
2D plate (variable mesh size) 7.83% 4.52% 2.98% 2.34% 2.16% 2.07%
@@ 5 Conclusion
Despite these advancements, our approach has its limitations. Primarily, the current model architecture is designed to generalize across various meshes of different size but does not extend this generalization to differing domain shapes. Furthermore, our evaluation focuses solely on the model’s capability in computing steady-state solutions. Further research is needed to extend PI-DCON to also incorporate dynamic responses, thereby enabling it to tackle time-dependent PDEs. Additionally, while our model currently relies on the basic multi-layer perceptron structure, incorporating more sophisticated architectures, such as Attention Mechanisms Vaswani et al. (2017), may further enhance its performance.

### v2

@@ 4.3 Comparison with data-driven neural operators
One of the main features of PI-DCON is that it is a data-free approach. It doesn’t require FE simulation runs as a training set. In this section, we seek to investigate its efficiency and accuracy in comparison with data-driven (supervised) training approaches. In doing so, we consider the data-driven DeepONet and also create a data-driven variant of our proposed DCON architecture for comparison. We first compare the performance of data-driven DCON with that of data-driven DeepONet data-driven version of Wang et al. (2022) in Table LABEL:table.data_driven_com. We observe that the data driven DCON model achieved 3.8% accuracy improvement for the Darcy flow problem and 20.7% accuracy improvement for the 2D plate problem, compared with the best baseline method. Also, when we compare our model’s performance on data with variable discrete representations to those with a fixed discrete representations, we see a small drop in the prediction accuracy, which is comparable to the drop observed in PI-DCON.
@@
Dataset $L_{2}$ Relative error
Data-driven DeepONet Data-driven Wang et al. (2022) Data-driven DCON
Darcy flow (fixed discrete representation) 3.74% $\pm$ 0.97% 1.82% $\pm$ 0.44% 1.66% $\pm$ 0.37%
2D plate (fixed discrete representation) 3.78% $\pm$ 1.01% 1.98% $\pm$ 0.41% 1.57% $\pm$ 0.31%
Darcy flow (variable discrete representation) - - 2.12% $\pm$ 0.49%
2D plate (variable discrete representation) - - 1.78% $\pm$ 0.34%
In order to demonstrate the advantage of physics-informed (data-free) training of DCON, in Figure 9 we show the required time needed for building models at different accuracy levels. The horizontal (dashed) line represents the required CPU time for generating training datasets for data-driven training of neural operators. In both darcy flow and 2d plate problems, the lowest error rate achieved by PI-DCON is smaller than that of data-driven DeepONet but greater than that of data-driven DCON. Also, PIDCON can obtain comparable results with the improved DeepONet architecture of Wang et al. (2022). This suggests that neural operators trained in a data-driven fashion tend to outperform those trained using physics-informed approaches when the same model architecture is used. However, it can be seen that PI-DCON convergence is superior compared to data-driven DeepONet and data-driven version of Wang et al. (2022). This shows the enhanced architectural design in DCON allows for faster convergence compared to the conventional DeepONet architectures.
@@ 4.4 Ablation studies
Darcy flow (fixed discrete representation) 10.10% 4.20% 2.50 % 2.15% 2.06% 2.03%
2D plate (fixed discrete representation) 5.92% 2.57% 1.77% 1.56% 1.43% 1.38%
Darcy flow (variable discrete representation) 12.58% 5.01% 3.42% 2.83% 2.66% 2.59%
2D plate (variable discrete representation) 7.83% 4.52% 2.98% 2.34% 2.16% 2.07%
We also seek to investigate our model performance with varying correlation lengths of the Gaussian Process and different amounts of training data. To do so, we train PI-DCON models using different numbers of training samples (specifically, 400, 300, 200, and 100 samples) for 500 epochs and report the prediction accuracy over 100 unseen samples with different correlation lengths $l$ in Figure  10. Our results show that increasing the number of training samples improves accuracy. For Gaussian processes with smaller correlation lengths, more training samples are required to achieve satisfactory performance. Additionally, if the training dataset size falls below a certain threshold, model performance significantly deteriorates. For instance, in the 2D plate problem, the relative error of the prediction increases from 5.66 to 15.22 when the dataset size decreases from 200 to 100, whereas it only increases from 3.42 to 5.66 when the dataset size decreases from 300 to 200.
(a) Darcy flow problem
(b) 2D plate problem
@@ 5 Conclusion
Despite these advancements, our approach has its limitations. Primarily, the current model architecture is designed to generalize across various discrete representations but does not extend this generalization to differing domain shapes. Furthermore, our evaluation focuses solely on the model’s capability in computing steady-state solutions. Further research is needed to extend PI-DCON to also incorporate dynamic responses, thereby enabling it to tackle time-dependent PDEs. Also, the evaluation on other well-studied PDE problems, such as Eikonal equation and Stokes equation, can also be investigated. Additionally, while our model currently relies on the basic multi-layer perceptron structure, incorporating more sophisticated architectures, such as Attention Mechanisms Vaswani et al. (2017), may further enhance its performance.
In addition to developing more innovative architectures, investigating new physics-informed training algorithms can also provide options for improvement. As pointed out by Wang et al. (2022), the balance between the PDE residual loss and the boundary condition loss significantly impacts model performance. Therefore, future works can focus on investigating various training algorithms that have shown promising results in PINN training, such as Variational PINNs (VPINNs) Kharazmi et al. (2019) and boundary distance function approximation Sukumar and Srivastava (2022).